An insurance market consists of high-risk patients, who average $40,000 in spending per year, and low-risk patients, who average $1,000 per year. Overall, low-risk patients represent 90 percent of the population. What would average spending be for a population like this?
Answer #1

Average spending = 0.10(40,000) +0.90(1,000) = 4,000 + 900 = $ 4,900